OpenStudy (anonymous):

The geometric mean of x-3 and x+4 is x. Find all possible values of x.

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

I also need to know:1.) the geometric mean of x and x^2 is 8. ???what are possible values of x??? 2.) the geometric mean of x+1 and 12x is 6x. x=?? 3.) the geometric mean of x-3 and 2x+8 is x+4...x=?? 4.) the geometric mean of x+1 and 3x+1 is 3x-1...x=??